# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ C.2 Raw Relationship between Income and Expenditures
  • [20] Tokyo Chamber of Commerce.Tokyoshinai shōshōkōgyō kinyūchōsa siryō (Survey report on the finance for small commercial and industrial enterprises in Tokyo City) [in Japanese]. Tokyo Chamber of Commerce, 1918.
@@
  • [21] Tokyo City Social Affairs Bureau.Tokyo shinai saimin no nyūshichi nikansuru chōsa (Survey on the use of pawnshops among the poor in Tokyo City) [in Japanese]. Tokyo City Social Affairs Bureau, 1921.
@@
  • [22] Tokyo City Office.Tokyo shi tōkei nenpyō, dai jyūnanakai (Annual Statistics of the City of Tokyo, vol.17) [in Japanese]. Tokyo: Tokyo City Office, 1921.
@@
  • [23] Tokyo City Office.Tokyo shi tōkeisho, dai jyūhachi kai (Annual Statistics of the City of Tokio, volume. 18). [in Japanese] Tokyo: Tokyo City Office, 1922b.
@@
  • [24] Tokyo City Office.Tokyo shi tōkeisho, dai jyūkyū kai (Annual Statistics of the City of Tokio, volume. 19). [in Japanese] Tokyo: Tokyo City Office, 1923.
@@
  • [25] Tokyo City Office. Tokyo shi shisei tōkei genpyō, dai ikkan (Tokyo City Municipal Statistics Original Table, vol.1) [in Japanese]. Tokyo: Tokyo City Office, 1922a.
@@
  • [26] Tokyo City Office. Chūshō shōkōgyōsha no kinyūmondai (Chūshō shōkōgyōsha shinkō chōsakai siryō, daini) (Financial issues for small and medium commercial and industrial enterprises (Material of the Committee for the Promotion of Small and Medium Commercial and Industrial Development, volume 2)) [in Japanese]. Tokyo: Tokyo City Office, 1935.
@@
  • [27] Tokyo City Statistics Division. Tokyo shi oyobi kinkō chōson rōdōtōkeigenpyō (dai ikkai, dai nikan) (Labor Statistics Table for Tokyo City and Suburbs, issue 2 of volume 1) [in Japanese]. Tokyo: Tokyo City Office, 1926a.
@@
  • [28] Tokyo City Statistics Division. Shichiyagyō no tōkei chōsa (Statistical Survey of the Pawnshop Business) [in Japanese]. Tokyo: Tokyo City Office, 1926b.
@@
  • [29] Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiissatsu toshikinyū no gaikyō (Survey of Financial Institutions among Common People, volume. 1, Overview of the Financial Institutions in the Cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925a.
@@
  • [30] Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, dainisatsu toshichochiku ginkō (Survey of Financial Institution among Common People, volume. 2, Municipal Savings Banks) [in Japanese]. Tokyo Institute for Municipal Research, 1925b.
@@
  • [31] Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daisansatsu shigaichi shinyōkumiai (Survey of Financial Institutions among Common People, volume. 2, Credit Unions in the cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925c.
@@
  • [32] Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, volume. 2, Pubic pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
@@
  • [33] Tokyo Legislative Research Association. Chokin to Mujin (Saving and Mutual Loan Association) [in Japanese]. Tokyo: Tokyo Legislative Research Association, 1915.
@@
  • [34] Umemura, Mataji, Takamatsu, Nobukiyo, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.13. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1983.
@@
  • [35] Umemura, Mataji, Akasaka, Keiko, Minami, Ryoshin, Takamatsu, Nobukiyo, Arai, Kurotake, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.2. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1988.
@@
  • [36] Water Works Association. Jyōsuidō tōkei oyobi hōkoku (Statistics and Reports of the Water Supply, Vol. 1). [in Japanese] Tokyo: Water Works Association, 1922.
@@
  • [1] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
  • [2] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
  • [3] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
  • [4] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
  • [5] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
  • [6] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
  • [7] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
  • [8] The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).

### 25 Feb 2024 v6

@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
  • (42) Tokyo Chamber of Commerce.Tokyoshinai shōshōkōgyō kinyūchōsa siryō (Survey report on the finance for small commercial and industrial enterprises in Tokyo City) [in Japanese]. Tokyo Chamber of Commerce, 1918.
  • (43)
  • (44) Tokyo Chamber of Commerce and Industry.Tokyoshinai niokeu kourigyōkeiei narabini kinyūchōsa (Survey on retail business management and finance in Toky City) [in Japanese]. Tokyo Chamber of Commerce and Industry, 1937.
  • (45)
@@
  • (46) Tokyo City Social Affairs Bureau.Tokyo shinai saimin no nyūshichi nikansuru chōsa (Survey on the use of pawnshops among the poor in Tokyo City) [in Japanese]. Tokyo City Social Affairs Bureau, 1921.
  • (47)
@@
  • (48) Tokyo City Office.Tokyo shi tōkei nenpyō, dai jyūnanakai (Annual Statistics of the City of Tokyo, vol.17) [in Japanese]. Tokyo: Tokyo City Office, 1921.
  • (49)
@@
  • (50) Tokyo City Office.Tokyo shi tōkeisho, dai jyūhachi kai (Annual Statistics of the City of Tokio, volume. 18). [in Japanese] Tokyo: Tokyo City Office, 1922b.
  • (51)
@@
  • (52) Tokyo City Office.Tokyo shi tōkeisho, dai jyūkyū kai (Annual Statistics of the City of Tokio, volume. 19). [in Japanese] Tokyo: Tokyo City Office, 1923.
  • (53)
@@
  • (54) Tokyo City Office. Tokyo shi shisei tōkei genpyō, dai ikkan (Tokyo City Municipal Statistics Original Table, vol.1) [in Japanese]. Tokyo: Tokyo City Office, 1922a.
  • (55)
@@
  • (56) Tokyo City Office. Chūshō shōkōgyōsha no kinyūmondai (Chūshō shōkōgyōsha shinkō chōsakai siryō, daini) (Financial issues for small and medium commercial and industrial enterprises (Material of the Committee for the Promotion of Small and Medium Commercial and Industrial Development, volume 2)) [in Japanese]. Tokyo: Tokyo City Office, 1935.
  • (57)
@@
  • (58) Tokyo City Statistics Division. Tokyo shi oyobi kinkō chōson rōdōtōkeigenpyō (dai ikkai, dai nikan) (Labor Statistics Table for Tokyo City and Suburbs, issue 2 of volume 1) [in Japanese]. Tokyo: Tokyo City Office, 1926a.
  • (59)
@@
  • (60) Tokyo City Statistics Division. Shichiyagyō no tōkei chōsa (Statistical Survey of the Pawnshop Business) [in Japanese]. Tokyo: Tokyo City Office, 1926b.
  • (61)
@@
  • (62) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiissatsu toshikinyū no gaikyō (Survey of Financial Institutions among Common People, volume. 1, Overview of the Financial Institutions in the Cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925a.
  • (63)
@@
  • (64) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, dainisatsu toshichochiku ginkō (Survey of Financial Institution among Common People, volume. 2, Municipal Savings Banks) [in Japanese]. Tokyo Institute for Municipal Research, 1925b.
  • (65)
@@
  • (66) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daisansatsu shigaichi shinyōkumiai (Survey of Financial Institutions among Common People, volume. 2, Credit Unions in the cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925c.
  • (67)
@@
  • (68) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, volume. 2, Pubic pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
  • (69)
@@
  • (70) Tokyo Legislative Research Association. Chokin to Mujin (Saving and Mutual Loan Association) [in Japanese]. Tokyo: Tokyo Legislative Research Association, 1915.
  • (71)
@@
  • (72) Umemura, Mataji, Takamatsu, Nobukiyo, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.13. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1983.
  • (73)
@@
  • (74) Umemura, Mataji, Akasaka, Keiko, Minami, Ryoshin, Takamatsu, Nobukiyo, Arai, Kurotake, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.2. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1988.
  • (75)
@@
  • (76) Water Works Association. Jyōsuidō tōkei oyobi hōkoku (Statistics and Reports of the Water Supply, Vol. 1). [in Japanese] Tokyo: Water Works Association, 1922.
  • (77)
@@
  • (1)
  • (2) The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
  • (3)
  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
  • (7)
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).
  • (12)